OpenStudy (anonymous):

Given sec s=5/3, where sin s<0, and cot t=1, where 0

OpenStudy (anonymous):

its long

OpenStudy (mertsj):

If the secant is 5/3 the cosine is 3/5. If the cot t is 1 and the angle is in the first quadrant, then the angle is 45 degrees and the sin and cos are both sqrt2/2

OpenStudy (mertsj):

sin(s-t)=sin s cos t - cos s sin t

OpenStudy (mertsj):

And cos(s+t)=cos s cos t - sin s sin t

OpenStudy (anonymous):

so what is s t=pie/4

OpenStudy (anonymous):

wha is s